OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. my answer is y=-1/4x-7/2

OpenStudy (wasiqss):

m=-1/4

OpenStudy (wasiqss):

sorry i got it wrong

OpenStudy (anonymous):

but i get this Y=-1/4x+b -3=-1/4(2) +b -3=-1/2+b b=-7/2 Y=-1/4x-7/2

OpenStudy (wasiqss):

you are right

OpenStudy (wasiqss):

good work

OpenStudy (anonymous):

thank u r the best :)

OpenStudy (wasiqss):

noooo i was right lol

OpenStudy (wasiqss):

yeah you got it wrong

OpenStudy (anonymous):

no i didnt

OpenStudy (wasiqss):

y=-1/4x-5/2

OpenStudy (anonymous):

how show me work

OpenStudy (wasiqss):

-3=-1/2+c c=-5/2

OpenStudy (wasiqss):

okay?

OpenStudy (wasiqss):

i should admire that your university algebra course is quite easy lol

OpenStudy (wasiqss):

if c=-5/2 m=-1/4 then y=mx+c y=(-1/4)x-5/2

OpenStudy (anonymous):

ok i get it

OpenStudy (wasiqss):

good ;)

OpenStudy (anonymous):

u add 1/2 not -

OpenStudy (wasiqss):

huh? but the final eq is right